Suppose V is finite-dimensional with dim V greater or equal to 2. Prove that there exist S, T €L(V, V) such that ST is not equal to T S


1
Expert's answer
2021-06-24T12:37:31-0400

solution:-

N1,N2 "\in" L(V,V) as follows

N1("\upsilon"1) =0, N1("\upsilon"2 )="\upsilon"(1)

N2("\upsilon"1)="\upsilon",N2("\upsilon"2)=0 (2)


Then for any vector w=a"\upsilon"1 + b"\upsilon"2 we have


N2N1(w)= aN2N1("\upsilon"1)+bN2N1("\upsilon"2)=b"\upsilon"(3)

but


N1N2(w)=aN1N2("\upsilon"1) + bN1N2("\upsilon"2)=a"\upsilon"1 (4)



we see from 3 and 4 the linear independence of "\upsilon"1,"\upsilon"2 that

N1N2(w)"\neq" N2N1(w)

unless a=b=0 that is ,unless w=0. Thus,

N1N2 "\neq" N2N1

as operators in L(V,V). In the event that dim V=n>2, we may build upon a construction of N1,N2follows : choosing a basis {"\upsilon"1,"\upsilon"2......,"\upsilon"n } for V, we define N1,N2 on "\upsilon"1,"\upsilon"2


as the above set

N1("\upsilon"i)=N2("\upsilon"i)=0

for 3"\leq"i"\leq"n, then for any w="\sum" ai"\upsilon"i"\in"V we have as above

N1N2(w) "\neq" N2N1(w)

providing at least one of a1,a2"\neq" 0, thus

N1N2"\neq" N2N1.